OpenStudy (anonymous):

If it takes half a man half a day to eat half a hen, how long does it take for a whole man to eat a whole hen?

OpenStudy (anonymous):

its a stupid question

OpenStudy (anonymous):

Half a day.

OpenStudy (anonymous):

a whole day , if anything is the answer it will be that

OpenStudy (anonymous):

You'd be surprised at the number of different answers I've gotten from people.

OpenStudy (anonymous):

yeh and they are all stupid

OpenStudy (anonymous):

Only two people gave half a day as the answer, including ag. I go along with the half a day answer.

OpenStudy (anonymous):

well then u are stupid :P

OpenStudy (anonymous):

how can "half a man" eat anything? by half a man they must mean a whole man

OpenStudy (anonymous):

I still think it'll be half a day. The amount and number have doubled, so time should remain the same. This is how I reasoned it out: 1/2M - 1/2 H - 1/2 D 1 M - 1/2 H - (1/2 * 1/2) = 1/4 D 1 M - 1H - (1/4 * 2) = 1/2 D

OpenStudy (anonymous):

If half a man is eating half the hen, what is the other half of him doing at the same time? Isn't the other half of him eating the other half of the hen?

OpenStudy (anonymous):

I am convinced that 1/2 day is the answer, too.

OpenStudy (anonymous):

whatever, I am not going to argue, I dont care if u get it right or wrong, either way the question is extremely badly worded. How can half a man eat ?. If anything it would be one person that is tired , define half a person eating? , If I were to chop you in half could you still eat?

OpenStudy (anonymous):

remember, maths is meant to be a bit bout logic thinking, not just pure memeorisation

OpenStudy (anonymous):

This is good. A math professor also stated that the question was badly worded.

OpenStudy (anonymous):

1day

OpenStudy (anonymous):

Iff the man were cut in two, then it would take him forever to eat a hen, because he'd be dead. If he remains whole and stays alive long enough to eat the hen, then both halves of the man are eating the hen at the same time.

OpenStudy (anonymous):

One man claimed the answer was 1/8 day, and he had a degree in math. Now that's incredible.

OpenStudy (anonymous):

His reasoning was that ya cube the deuce and divide into the ace to get an eighth of a day.

OpenStudy (anonymous):

It isn't about mugging, it's just that quirky thing about maths. In real life, you can't have a man cut in half, but in maths, it's perfectly reasonable to assume that. If you're uncomfortable with those numbers, we could just change the question by doubling everything, because it's all proportional. Having said that, if a man eats a hen in a day, 1 man will eat 2 hens in one day because each man would eat a hen in a day. To get the answer to this sum, just divide all numbers in the above reasoning by 2, and you get your answer. I hope that sounds more logical, but honestly, the numbers don't matter. It's only maintaining the ratio that does.

OpenStudy (anonymous):

In order to carry the analysis to two men and two hens, we should assume both eat at the same rate and that the hens are identical, since, with the original data, one man eats eats at exactly the same rate as himself, and the hen is identical to itself.

OpenStudy (anonymous):

1/2 man eats 1/2 hen in 1/2 day 1/2 man eats 1 hen in 1 day 1 man eats 2 hens in 1 day 1 man eats 1 hen in 1/2 day

OpenStudy (anonymous):

We should also assume that both men start and finish eating at the same time, winde one man eats at the same time as himself.

OpenStudy (anonymous):

I also agree with 1/2 day.

OpenStudy (anonymous):

A very important assumption is that the men live long enough to finish eating the hens.

OpenStudy (anonymous):

question needs clarification. what is half a man? get the semantic ambiguities out of the way before you solve it.